OpenStudy (anonymous):

Given the slope and one point, write the equation of the line in slope intercept form m=3,(-3,5)

OpenStudy (anonymous):

Equation of a line, y = mx+b form.

OpenStudy (anonymous):

i have the answer. y-3x+14. can you show me how to get the answer. i have to do my eca's next week

OpenStudy (anonymous):

Just use point - slope form \[\huge y-y _{1}=m(x-x _{1})\] Then transform it into y = mx+b Where m is the slope, x1 and y1 are your coordinates.

OpenStudy (anonymous):

can you help me how to plug the numbers

OpenStudy (anonymous):

What do you mean, you should be able to do that.

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you. i do my classes online and i did this part of algebra like 6 months ago and i forgot

OpenStudy (anonymous):

so i got this i don't what number to put in for m